# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

### v1

@@ 2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. Alternatively, parts-based approaches like ProtoPNet (Chen et al., 2019; Donnelly et al., 2021) learn a predefined number of prototypical parts per class. Each class-specific prototype is an image patch represented by a $1\times 1$ spatial unit in the latent space. The network classifies images by summarizing the similarities of all the latent image patches to all prototypes using a fully connected layer. PIPNet (Nauta et al., 2023) aligns learned prototypes with human perception by optimizing the model to assign the same prototype to two different views of the same augmented image patch. ProtoPShare (Rymarczyk et al., 2020) merges similar prototypes after training is complete, while ProtoPool (Rymarczyk et al., 2022) directly learns class-shared prototypes. Other approaches additionally replace the linear output layer with other interpretable models. ProtoKNN (Ukai et al., 2023) classifies the similarity scores using a k-nearest neighbor (KNN) and ProtoTree Nauta et al. (2021) learn a decision tree.
HyperPgNet learns a predefined number of prototypes per concept, not per class. These concepts are provided as pixel-level annotations extracted by a new labeling pipeline based on the foundation models DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024). Similar to part-based prototypes, the concept prototypes span a $1\times 1$ spatial unit in the latent space. The prototypes are exclusive to each concept but shared among the classes.
Prototype learning is also applied to image segmentation. ProtoSeg (Sacha et al., 2023) uses prototypical parts a latent space with high spatial resolution for image segmentation. Another method proposes non-learnable prototypes (Zhou et al., 2022), which are trained by clustering in the latent space instead of being optimized via Backpropagation. ProtoGMM (Moradinasab et al., 2024) builds on this by updating class-specific prototypes like a Gaussian Mixture Model (GMM) using an Expectation-Maximization (EM) algorithm (Liang et al., 2022) with fixed mixture weights. MGProto (Wang et al., 2023) also learns mixture weights for their prototype GMM model.
HyperPgNet employs HyperPg, a novel prototype representation using Gaussian distributions on a hypersphere’s surface with learned parameters: direction $\bm{{\alpha}}$, mean $\mu$ and standard deviation (std) $\sigma$. Unlike ProtoGMM, which learns a multivariate Gaussian distribution in Euclidian space, each HyperPg prototype is defined by a directional vector $\bm{{\alpha}}$ and a $1D$-Gaussian distribution of cosine similarity values with mean $\mu$ and standard-deviation $\sigma$ around this vector. The use of cosine similarity for prototype learning has been shown to improve classification (Mettes et al., 2019) and has been widely adopted. (e.g., Donnelly et al., 2021; Rymarczyk et al., 2022; Sacha et al., 2023). HyperPgNet advances this by computing prototype activations as the probability density of each HyperPg prototype’s Gaussian distribution, allowing each prototype to adapt to the spread of the latent cluster.
@@ 3 Prototypical Gaussians on the Hypersphere
Figure 1: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. $L_{2}$ prototypes compute the Euclidian distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidian space and compute a probability density function (PDF). HyperPg prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@ 3.3 Gaussian Prototypes on the Hypersphere - HyperPg
Prototypical Gaussians on the Hypersphere (HyperPg) combine the advantages of Gaussian and hyperspherical prototypes. HyperPg prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPg prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPg’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
@@
(a) $\mu=1$
(b) $\mu=0.6$
(c) $\mu=0$
(d) $\mu=-1$
Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$. The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$. The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution. For $\mu=1$, the distribution is aligned with the vector; for $\mu=-1$ it is on the opposite side of the hypersphere. Due to the cosine similarity, setting $11>\mu>-$ 1 interpolates the area of highest probability density between both poles, leading to a ring shape on the hypersphere’s surface.
@@
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation. This activation pattern for the cosine similarity would typically require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPg achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes, as a single HyperPg prototype can cover multiple clusters simultaneously.
@@ 4.1 Prototype Losses
$L_{\mathrm{Den}}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
@@ 4.3 Network Architecture
Fig. 3 illustrates HyperPgNet’s Architecture for interpretable image classification. HyperPgNet uses a pretrained, off-the-shelf feature encoder as a backbone, such as the convolution pyramid from a CNN like ConvNeXt-tiny (Liu et al., 2022) or a transformer architecture like Segformer (Xie et al., 2021). The high-dimensional feature maps provided by the backbone are passed through a neck, consisting of two $1\times 1$ convolution layers with ReLU activation in between, projecting into a smaller latent space with $256$ channels, followed by Layer Normalization. The hyperparameter of 256 channels aligns with previous work (e.g., Rymarczyk et al., 2022).
HyperPg’s similarity function computes the truncated Gaussian PDF over the cosine similarity of the latent feature maps with the HyperPg prototype anchors $\bm{{\alpha}}$. HyperPgNet implements the prototype learning part of the network as a module with two layers: First, a prototype learning layer computes the cosine similarity of learnable HyperPg prototype anchors $\alpha$ to the latent feature map $z$, similar to other prototype learning architectures. Second, the new Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the truncated Gaussian PDF over the activations of the previous layer.
This implementation of the HyperPg Module with two internal layers supports swapping out the prototype similarity measure to other measures in the future. Additionally, the Density Estimation Layer concept could be applied to other architectures to learn a Gaussian distribution over the previous layer’s activations. Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
HyperPgNet is trained for image classification with a multi-objective loss function. This loss function is defined as
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Den}}L_{\mathrm{Den}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
where $L_{\mathrm{CE}}$ is the Crossentropy loss over network predictions and ground truth image labels. By choosing different values for $\lambda$, the model’s tradeoff between task performance and interpretability can be adjusted.
@@ 5 Concept Extraction at Scale
The CUB-200-2011 dataset (Wah et al., 2011) contains 11,788 images of birds across 200 sub-species. Each image is annotated with up to 15 part locations, such as head, tail and back. The dataset includes 312 binary attributes describing patterns and colors of these parts, or additional shape information like “duck-like” or “owl-like”. For concept learning, 30 attributes are manually selected, excluding small concepts like eye-color or class exclusive attributes like “duck-like”.
To extract pixel level annotation masks, the point annotations for the attributes are given to a Segment Anything 2 (SAM2) model(Ravi et al., 2024). Using the point locations belonging to other attributes as negative points, segmentation masks are generated (see Fig. 4). The segmentation mask is then assigned the attribute from the dataset annotations to define a concept.
@@
The Stanford Cars dataset (Krause et al., 2013) does not provide part annotations. This dataset contains 16,185 images of 196 car classes, categorized by make, model and year (e.g., BMW M3 coupe 2012). The dataset provides only bounding boxes for the main subject matter, useful for foreground-background separation. To obtain pixel-level annotations, a list of 10 car parts (e.g., “wheel”, “headlight”, “radiator”) is defined. A Grounding DINO model (Liu et al., 2023) uses this list to generate bounding boxes for parts in the image. A post-processing step removes possible outliers by eliminating bounding boxes outside the original bounding box provided by the dataset.Finally, SAM2 generates pixel-level segmentation masks (see Fig. 4). For Stanford Cars, only the part label is used as the concept label as the dataset does not provide additional part attributes. The automated pipeline labeled the entire Stanford Cars dataset within 2h on an NVidia 4060ti with 16GB VRAM and can be easily adapted to other datasets.
@@ 6 Experiments
The experiments were performed on two datasets: CUB-200-2011 bird species classification dataset (Wah et al., 2011) and the Stanford Cars dataset (Krause et al., 2013). Two variants of HyperPgNet were tested, with and without the $L_{\mathrm{RRC}}$ loss. HyperPgNet was compared to ProtoPNet with both standard $L_{2}$ and the newly introduced HyperPg prototypes, and Concept Bottleneck Models (CBM) with joint training scheme (Koh et al., 2020). Segformer and ConvNext served as Transformer and CNN baselines, respectively. All models were implemented with pretrained Segformer backbones in Pytorch. However, HyperPgNet’s performance is independent of the chosen backbone architecture (see Appendix B.1).
In contrast to prior work, the training and testing was done on full images, without crops to the bounding boxes provided by the datasets. For data augmentation RandomPerspective, RandomHorizontalFlip, and RandomAffine were used in an online manner. Thus, for CUB-200-2011, each class has 30 training images, instead of the offline augmentation to 1200 images used in prior work (e.g., Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023). All models were trained with batch size 96 until convergence. The experiments were performed on a workstation with a single NVidia RTX 3090 GPU (24GB VRAM) per model.
@@ 6.1 Quantitative Results
Table 1: Test Top-1 Accuracy with Segformer Backbone. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
CUB-200-2011 Stanford Cars
# Prototypes Accuracy [%] # Prototypes Accuracy [%]
BB Segformer Baseline - 17.7 - 1.9
ConvNeXt Baseline - 74.2 - 57.5
CBM - 75.7 - 79.9
PP ProtoPNet 2000 68.0 1960 86.4
ProtoPNet + HyperPg 2000 70.5 1960 87.4
CAP HyperPgNet - $L_{\mathrm{RRC}}$ 300 76.5 180 88.6
HyperPgNet + $L_{\mathrm{RRC}}$ 300 74.1 180 81.2
Figure 5: Test Accuracy with Segformer Backbone.
Table 1 shows the top-1 accuracy for models on the CUB-200-2011 (Birds) and Stanford Cars (Cars) datasets using the Segformer backbone. The Test Accuracy per Epoch curves are visualized in Fig. 5. The Segformer baseline overfit to the small training datasets (17.7% and 1.9% test accuracy), while the ConvNeXt baseline performed better (74.2% and 57.5%). CBM achieved high results quickly, scoring 75.7% on Birds and 79.9% on Cars. ProtoPNet required the most training time, converging only after about 490 epochs, scoring 68% on Birds and 86.4% on Cars. Switching ProtoPNet’s prototypes from $L_{2}$ to HyperPg improved both training speed and accuracy, converging in 200 epochs with 70.5% and 87.4% test accuracy.
HyperPgNet further improved on this by learning human-defined concepts. HyperPgNet with concept aligned prototypes, but without $L_{\mathrm{RRC}}$, outperformed the other models. It achieved 76.5% accuracy on Birds and 88.6% on Cars after only 40 training epochs. With the “Right for the Right Concept” loss, the model retained the high learning speed and performance, although the test accuracy dropped slightly to 74.1% and 81.2%. This performance loss is offset by increased transparency due to more precise concept prototypes. For additional ablation studies, see Appendix B.
6.2 Qualitative Results
HyperPg prototypes do not only enhance the models learning speed and task performance, but also the interpretabilty. Prior work built activations by overlaying the up-sampled latent space activation over the image. Inspired by GradCAM, Fig. 6 shows the gradient of the prototype activations backpropagated to the input image. Class-based prototypes focus on regions relevant for the prediction, but do not follow human defined concepts. With concept-alignment, the prototype focuses more towards the white head of the bird. But as the concept alignment is only encouraged by $L_{\mathrm{Clst}}$ and $L_{\mathrm{Sep}}$ the prototype still activates on other bright areas on the birds body. With the addition of $L_{\mathrm{RRC}}$ to the training regime, the prototypes are forced to focus mainly on the annotated region. Thus, the head prototype ignores the wing area, which is learned by a different prototype. The overall learned concept prototypes fit more precisely to the images’ annotated areas. For additional examples and latent space analyses, please refer to Appendix C.
(a) ProtoPNet + HyperPg
(b) HyperPgNet - $L_{\mathrm{RRC}}$
(c) HyperPgNet + $L_{\mathrm{RRC}}$
Figure 6: Prototype Gradient Maps for different HyperPgNet models.

### v2

@@ 2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. ProtoPNet (Chen et al., 2019) introduced the idea of prototypical parts. In this setting, each prototype is a latent patch of the input image, commonly a $1\times 1$ latent patch. The prototypes are each associated with a single class and learned via backpropagation without additional information.
HyperPgNet builds on the idea of prototypical parts, but learns the prototypes in a transparent, concept-aligned manner. The prototypes are not class exclusive, but class shared. Each prototype corresponds to a human defined concept. To provide the required concept annotations, we propose a labeling pipeline based on foundation models such as DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024). HyperPgNet learns the concept-aligned prototypes by only adapting the loss functions and providing additional annotations, without requiring changes to the feature encoder like PIPNet (Nauta & Seifert, 2023) or Lucid-PPN (Pach et al., 2024), which uses a hybrid input head to disentangle prototype color and shape.
Multiple successors build on the idea of ProtoPNet. ProtoPShare (Rymarczyk et al., 2020) and ProtoPool (Rymarczyk et al., 2022) use additional optimizations to learn class-shared prototypical parts. Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement. Other work replace the linear output layer of ProtoPNet with other interpretable models: ProtoKNN (Ukai et al., 2023) uses a k-nearest neighbor classifier and ProtoTree (Nauta et al., 2021) employs a decision tree.
All these models change how the prototype activations are further processed. However, they all rely on the point-based prototype formulation introduced by ProtoPNet based on the $L_{2}$ or cosine similarity. HyperPgNet proposes a novel prototype formulation based on Gaussian distributions on the hypersphere, without changing the downstream processing. This probabilistic view with learned mean and variance allows HyperPgNet to learn prototypes with different degrees of specialization. Furthermore, the statistical confidence derived from the learned variance could be integrate in further downstream processing.
The view of prototype learning as clustering in latent space was introduced for image segmentation (Zhou et al., 2022). The image patches are points in latent space with the prototypes as cluster centers. ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2023) build on this idea to model the prototypes as Gaussian mixture models for image segmentation.
HyperPg adapts the idea of Gaussian prototypes used in image segmentation to hyperspherical latent spaces based on the cosine similarity. The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019) and also been applied in recent prototype learning works (Zhou et al., 2022; Ukai et al., 2023). This makes HyperPgNet the first model to employ probabilistic prototypes with learnable mean and variance in a hyperspherical space.
@@ 3 Prototypical Gaussians on the Hypersphere
Figure 1: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. $L_{2}$ prototypes compute the Euclidean distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF). HyperPg prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@ 3.3 Gaussian Prototypes on the Hypersphere - HyperPg
Prototypical Gaussians on the Hypersphere (HyperPg) combine the advantages of Gaussian and hyperspherical prototypes. HyperPg prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPg prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPg’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
@@
Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$. The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$ for 10k random samples. The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution. For $\mu=1$, the distribution is aligned with the vector; for $\mu=-1$ it is on the opposite side of the hypersphere. Due to the cosine similarity, setting $11>\mu>-$ 1 interpolates the area of highest probability density between both poles, leading to a ring shape on the hypersphere’s surface.
@@
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation. This activation pattern for the cosine similarity would typically require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPg achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes and is a major difference to the von Mises-Fisher distribution (see also Appendix A).
@@ 4.1 Prototype Losses
$L_{Density}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
@@ 4.2 Right Concept Loss
4.3 Multi-Objective Loss Function
To train a prototype learning network like HyperPgNet for downstream tasks like image classification, a multi-objective loss function is employed. This multi-objective loss function is defined as
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels. ProtoPNet uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ (Chen et al., 2019). For HyperPgNet, the different loss terms are weighted equally, meaning all $\lambda=1$.
@@ 4.4 Network Architecture
Fig. 3 illustrates HyperPgNet’s Architecture for interpretable image classification. HyperPgNet uses a pretrained feature encoder such as ConvNext-tiny (Liu et al., 2022) or MiT-B4 (Xie et al., 2021) as a backbone model. A neck consisting of two $1\times 1$ convolution layers with ReLU activation in between projects the high dimensional feature map of the backbone into a lower dimensional feature space.
The HyperPg prototype module is implemented with two layers: First, a prototype learning layer computes the cosine similarity of the learnable HyperPg anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck. Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer. Future extensions can adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
@@ 5 Concept Extraction at Scale
The CUB-200-2011 (Wah et al., 2011) annotates each image with up to 15 part locations (e.g., head, tail, wing). Each part is labeled with some attribute describing the coloring, pattern or shape. A Segment Anything 2 (SAM2) model (Ravi et al., 2024) creates pixel-level segmentation masks based on the part locations.
The Stanford Cars dataset (Krause et al., 2013) does not provide part level annotations. We defined a list of 10 car parts (e.g., wheel, headlight, radiator) and use a Grounding DINO model (Liu et al., 2023) to generate bounding boxes for each part. Finally, SAM2 creates pixel-level segmentation masks based on the bounding boxes. This automated pipeline labeled the 16k images of Stanford Cars within 2h on a NVIDIA 4060Ti with 16GB VRAM, and can be easily adapted to other datasets. Fig. 4 presents some examples for CUB-200-2011 and Stanford Cars.
@@ 6 Experiments
The experiments were performed on two datasets: CUB-200-2011 (CUB) with 200 bird species Wah et al. (2011) and Stanford Cars (CARS) with 196 car models Krause et al. (2013). Implementation details such as data preprocessing or model implementation are provided in Appendix B.
Birds Cars
# Prototypes MiT-B4 ConvNeXt # Prototypes MiT-B4 ConvNeXt
BB Baseline - 17.7 74.2 - 1.9 57.5
CBM - 75.7 77.9 - 79.9 81.0
PP ProtoPNet (PPN) 2000 68.0 68.1 1960 86.4 87.0
PPN + HyperPG 2000 70.5 65.0 1960 87.4 77.3
CAP HyperPgNet $-L_{\mathrm{RRC}}$ 300 76.5 76.9 180 88.6 88.9
(Ours) HyperPgNet $+L_{\mathrm{RRC}}$ 300 74.1 71.5 180 81.2 86.0
Table 1: Test Top-1 Accuracy. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
@@ 6.1 Quantitative Results
Figure 5: CUB Test Accuracy per Epoch.
Table 1 reports the top-1 accuracy for the tested models on the CUB and CARS datasets with different backbones. HyperPgNet achieves comparable results with the black box CBM and baseline models. Notably, HyperPgNet outperforms CBM on CARS, as it can differentiate between different modes for the same concept by learning multiple prototypes per concept. CBM is restricted to a binary concept presence prediction. Including ${\mathcal{L}}_{\mathrm{RRC}}$ slightly lowers HyperPgNet’s performance but is still comparable to the black-box models.
Fig. 5 shows the test accuracy over epochs for CUB with ConvNeXt backbone. The concept aligned models CBM and HyperPgNet are the fastest to converge at around 100 epochs. ProtoPNet requires over 400 training epochs until the best performance level is reached. However, switching from the standard $L_{2}$ based prototypes to HyperPg prototypes effectively halves the required training epochs.
6.2 Qualitative Comparison
Fig. 7 shows the pixel attribution of the highest activated prototype on each image for the tested prototype learning models with ConvNeXt backbone. In contrast to prior work (e.g., Chen et al., 2019; Ukai et al., 2023), this visualization is based on the prototype gradients, similar to saliency maps such as GradCAM (Selvaraju et al., 2017), which should avoid spatial misalignment often observed in prototype explanations (Sacha et al., 2024). The concept aligned prototypes learned by HyperPgNet focus more on the relevant parts of the image subject than the class-based prototypes of ProtoPNet. The network seems to focus correctly on either the overall color of the bird and ignoring the background (first three examples) or on specific markings (last three examples). The addition of the ${\mathcal{L}}_{\mathrm{RRC}}$ loss seems to further increase the main focus of the prototype activation on the relevant image regions. For ProtoPNet, no qualitative difference can be detected between the standard $L_{2}$ or HyperPg prototypes. With both prototype formulations, ProtoPNet is likely to also activate on parts of the image background. This is a fundamental weakness of training ProtoPNet in an opaque manner without alignment to relevant concepts, which HyperPgNet addresses.
On the CARS dataset, the effect of ${\mathcal{L}}_{\mathrm{RRC}}$ is more apparent. The defined concepts for this dataset are limited to car parts without fine-grained differentiation between patterns or shapes, such as the concepts on CUB. For this dataset, one of the most predictive image regions across models seems to be the front bumper of the cars, which has not been annotated with its own concept. As shown in Fig. 6, the concept-aligned prototypes still activate in this region. However, by training the model with the ${\mathcal{L}}_{\mathrm{RRC}}$ loss, the prototypes focus more on the annotated image regions.
Figure 6: Concept prototype activation after training with and without ${\mathcal{L}}_{\mathrm{RRC}}$ with MiT-B4 backbone.
Figure 7: Highest prototype activation per image with ConvNeXt backbone.
6.3 Ablation
# Prototypes CUB Accuracy
CBM 71.6
1 70.5
5 68.9
10 68.0
20 67.3
Table 2: Test Top-1 Accuracy on CUB with ResNet50 backbone with varying numbers of prototypes per concept.
We conduct an ablation on the number of prototypes per concept. Table 2 presents the test accuracies on CUB for HyperPgNet with ResNet50 backbone. The difference in task performance over the different model configurations is negligible. However, looking at some qualitative examples as depicted in Fig. 8 a difference in prototype quality can be observed. With fewer prototypes per concept, the model is more likely to also react to parts of the mage background in order to still achieve a low training error. With too many prototypes per concept, the latent space becomes too segmented. The prototypes start to focus on smaller subparts of the image subject, potentially decreasing the interpretability. The chosen configuration of 10 prototypes per concept seems to be a sweet spot regarding predictive performance and prototype quality for the experiments.
Figure 8: Qualitative Results for HyperPgNet with ResNet50 backbone and varying numbers of prototypes per concept.